• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name markowitz

    Log entry time 08:11:15 on April13,2001

    Entry number 60336

    keyword=shift summary

    00:00 Run 3481 underway at shift start (kin 30). Finished at 00:17

    00:17 started run 3482. finished at 00:50

    00:50 started run 3483; stopped at 00:52


    00:52 MCC started the QE measurement.


    01:04 beam back; started run 3484. Stopped at 01:40. 1158 kaons so far for kin 30.

    01:40 started run 3485. 02:01 finished. Final tally at kin 30: 1260 kaons

    BOILING STUDY: during study, always waited 2-3 minutes after current was set before starting run. Required no beam trips during a successful run. If a trip occurred at end of run, since we do not acquire any events, run still ok but refuse to take more events in same run. Changed current in about 10 uA steps.

    02:05 started run 3487 (target boiling at 92 uA). Stopped at 02:11 with 200K events

    02:17 started 3488, 82 uA. Stopped at 02:27, 270K events

    02:32 started 3489, 73 uA. Stopped at 02:36 when beam tripped off. 86K events.

    02:40 started 3490, 72 uA. Stopped at 02:51, 250K events

    02:55 started 3491, 62 uA. Stopped at 02:57 when beam tripped off. 18K events.

    03:00 started 3492, 62uA. Stopped at 03:16, 290K events

    03:21 started 3493, 52 uA. Stopped at 03:42, 308K events

    03:48 started 3494, 43 uA. Stopped at 04:06 when beam triped off. 177K events

    04:10 started 3495, 43 uA. Stopped at 04:18 when beam tripped off. 76K events

    04:25 started 3496, 34 uA. Stopped at 04:36 when beam tripped off. 67K events

    04:39 started 3497, 34 uA. Stopped at 05:09, 208K events


    moving to multi-foil carbon target.
    Changing prescales from 900/5000/60000/10000/1 to 1/400/50/550/1.


    05:30 started run 3498, 12C multi-foil at kin 30. Only see half of the foils (1, 5-9 but not 2-4).

    run 3499 we moved to y=-2 mm, x=0.1 mm. see more of the foils ubt only a hint of 2-4

    run 3500 moved beam to x=+2mm, y=0 and see all 9 foils nicely. Conclude that the target shift of 4 mm was real and for the 12C multi-foil is actually 5+ mm making the foils JUST on or off the beam without adjustment. So we move beam to +2mm and hit the targets again.


    05:50 Request an access to move the spectrometers. Pete, Rusty enter the hall.
    moved spectrometer from 12.5 to 21.18 degrees run 3502 is kin 29 angles

    07:00 moved to kin27 angles from upstairs and ran multifoil. Run 3503. ps at 1/60/25/700/1.


    07:30 angles set to kin 26, run 3504.

    07:45 angles set to kin24, run 3505.